OpenStudy (anonymous):

If r(x)=54x^6-12x^4-27x^3 and t(x)=-9x^3 find a.) (r/t)(x) b.)(r/t)(3) c.)(r/t)(-1) hope to get some answers and please explain

OpenStudy (anonymous):

take the derivative of r(x) with respect to x dr/dx next take the derivative of dt/dx use the chain rule dr/dt = dr/dx * dx/dt that gives you dr/dt

OpenStudy (anonymous):

Sorry that makes no sense to me

OpenStudy (anonymous):

I think he misread the questions. Assuming I AM reading it right (granted, the notation is a bit weird): For the first one, just divide r(x) by t(x) (divide each term in r(x) by (-9x^3). the other two are the same, but with a specific value of x to plug in.

OpenStudy (anonymous):

use long division to divide the polynomials

OpenStudy (anonymous):

I guess you could call it long division, but it's really just: \[\frac{54x^6-12x^4-27x^3}{-9x^3} = \frac{54x^6}{-9x^3} + \frac{-12x^4}{-9x^3} + \frac{-27x^3}{-9x^3} = ... \]

OpenStudy (anonymous):

I like that better so how do i apply it to the problem

OpenStudy (anonymous):

(Again, this assumed the notation is as I think) Just find that answer to that (the simplified form). Then let x = 3 and -1 for the second and third parts, respectively.

OpenStudy (anonymous):

so how do i apply it to (r/t)(x)

OpenStudy (anonymous):

... That is what (I assume) (r/t)x means...r(x)/r(t)... does it not?

OpenStudy (anonymous):

so my answer to (r/t)(x) is: -6^3+4/3x+3

OpenStudy (anonymous):

I believe so.

OpenStudy (anonymous):

ok i think i got it now